# A New Family of Binary Sequences via Elliptic Function Fields over Finite Fields of Odd Characteristics

### v1

@@ I-A Our Main Results and Techniques
where $\lfloor x\rfloor$ stands for the integer part of $x\in\mathbb{R}$ and correlation upper bounded by
$(2d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+2d.$
By taking explicit cyclic elliptic function fields, we extend the families of the binary sequences with new parameters, see Table I.
Now we overview the techniques which shall be used throughout this paper. The automorphism group of an elliptic function field $E$ over $\mathbb{F}_{q}$ is in the form of the semi-product $T_{E}\rtimes\mathrm{Aut}(E,O)$, where $\mathrm{Aut}(E,O)$ consists of the $\mathbb{F}_{q}$-automorphisms of $E$ fixing the infinity place $O$, and $T_{E}$ corresponds to the translation group $E$. Let $Q$ be a place of $E$ with $\deg(Q)=d\geq 2$ such that $\gcd(d,q+1+t)=1$ and by the Riemann-Roch theorem, we have $\ell(Q)=\dim(\mathcal{L}(Q))=\deg(Q)+1-g(E)=\deg(Q)=d$ because of $\deg(Q)=d\geq 2g(E)-1=1$. Obviously, we have the decomposition $\mathcal{L}(Q)=\mathbb{F}_{q}\oplus V$. The vector space $V$ is called the complementary dual space of $\mathbb{F}_{q}$ in $\mathcal{L}(Q)$. The elements in $V\setminus\{0\}$ can be listed as $\{z_{1},z_{2},\cdots,z_{q^{d-1}-1}\}$.
@@
Then the family of binary sequences $\mathbf{s}_{i}$ has length $q+1+t$ and size $q^{d-1}-1$. The problems of computing the correlation and linear complexity are transformed into counting the number of rational places of certain Kummer extensions $E_{j,k}=E(y)$ where $y^{2}=z_{j}\cdot\tau^{-u}(z_{k})$ for $1\leq j,k\leq q^{d-1}-1,0\leq u\leq q+t$ by the cyclic structure of $T_{E}$.
@@
Although our bound applies to more general cases of the finite fields, both the zero places and pole places of $z_{j}\cdot\tau^{-u}(z_{k})$ are ramified in the Kummer extensions while only the pole places of $z_{j}+\tau^{-u}(z_{k})$ are ramified in the Artin-Schreier extensions for $1\leq j,k\leq q^{d-1}-1,0\leq u\leq q+t$, which induce the differences between the two bounds. But the two bounds are closed and the actual correlation is much smaller than the theoretical bound as is shown in the numerical experiments.
For cryptographic applications, we explored the property of balance. Also by counting the rational places over the Kummer extensions, we have found that the theoretical value of the balance is upper bounded by
$(d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+d.$
By utilizing the software SageMath [2], it is also shown in table II that the actual difference is also much smaller than this theoretical bound.
@@ I-B Organization of This Paper
The rest of this paper is organized as follows: In Section II, we provide preliminaries on binary sequences, extension theory of function fields, elliptic function fields and Kummer extensions of the function fields. In Section III, we present a theoretical construction of binary sequences with low correlation, favorable balance and large linear complexity via cyclic elliptic function fields over finite fields with odd characteristic by utilizing the theory of the Kummer extensions and we also present some examples to specify our construction. In Section IV, we give a conclusion and some possible future work on the constructions of binary sequences.
@@ III-B The Balance and Correlation
Now we define $\Delta(\mathcal{S})\triangleq\max\{|N^{\prime\prime}_{1}(\mathbf{s})-N^{\prime\prime}_{0}(\mathbf{s})|,\ \mathbf{s}\in\mathcal{S}\}$ and $\Delta(\mathcal{S})$ is called the balance. In this subsection, we study the property of balance and it turns out that the problem is transformed into the analysis of the number of rational places for the Kummer extensions.
@@
$\begin{split}\Delta(\mathcal{S})&=\left|2N_{1}^{\prime\prime}(\mathcal{S})-(q+1+t)\right|\\ &\leq(d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+d.\end{split}$ (2)
where $\Delta(\mathcal{S})=\max\{|N^{\prime\prime}_{1}(\mathbf{s}_{i})-N^{\prime\prime}_{0}(\mathbf{s}_{i})|,1\leq i\leq q^{d-1}-1\}$.
@@
On the other hand, we have $N(E^{\prime}_{i})=2N^{\prime\prime}_{0}(\mathbf{s}_{i})-|\mathcal{T}(z_{i})|$ for $1\leq i\leq q^{d-1}-1$. From the Serre bound, we have the estimation for the number of rational places over the curve $E^{\prime}_{i}$ and it is given by
@@
Let $n$ be an odd positive integer, $q=3^{n}$ and $t=\pm 3^{(n+1)/2}$. Then there exists a family of binary sequences with length $q+1+t$, size q1q-1 and the correlation upper-bounded by $\mathrm{Cor}(\mathcal{S})\leq 5\cdot\lfloor 2\cdot 3^{n/2}\rfloor+3^{(n+1)/2}+4$.
@@
For $t=\pm 3^{(n+1)/2}$, there is a cyclic elliptic function field with $q+1+t$ rational places according to Lemma II.4. By Lemma II.5, there exists a place $Q$ in $E^{T_{E}}$ of degree $\deg(Q)=2$ that splits completely in $E$ as $\gcd(2,3^{n}+1+t)=1$. Thus, the corollary follows from Theorem III.6. ∎
@@
Let $q=5^{n}$ for a positive integer $n$. Let $t=\pm 5^{n/2}$ and $n$ be even. Then there exists a family of binary sequences with length $q+1+t$, size q1q-1 and correlation upper bounded by $\mathrm{Cor}(\mathcal{S})\leq 5\cdot\lfloor 2\cdot 5^{n}\rfloor+|t|+4=11\cdot 5^{n/2}+4$.
@@
For $t=\pm 5^{n/2}$, there is a cyclic elliptic function field with $q+1+t$ rational places according to Lemma II.4. By Lemma II.5, there exists a place $Q$ in $E^{T_{E}}$ of degree $\deg(Q)=2$ that splits completely in $E$ for the fact that $\gcd(2,5^{n}+1+t)=1$. Thus, the corollary follows from Theorem III.6. ∎
@@ III-D Numerical Results
Field Size Family Size $N^{\prime\prime}_{0}(\mathbf{s})$ $N^{\prime\prime}_{1}(\mathbf{s})$ $|N_{1}^{\prime\prime}(\mathbf{s})-N_{0}^{\prime\prime}(\mathbf{s})|$
81 80 4063 4037 26
243 242 5991 6159 168
729 728 18302 18148 154
Now we consider the actual correlation. Also let the field size be $q=3^{n}$. By the results in Theorem I, we can construct a family of binary sequences with size $3^{n}-1$, length $q$ and correlation upper bounded by $5\cdot\lfloor 2\sqrt{q}\rfloor+4$. We list some results calculated by SageMath [2] for this case in Table III.
@@
Field Size Sequence Length Family Size Theoretical Bound of Correlation Actual Correlation
@@
Field Size Sequence Length Family Size Theoretical Bound of Correlation Actual Correlation
@@ IV Conclusions and Future Work
In this paper, we extend the construction of binary sequences given by Jin et al. in [9] via the cyclic elliptic function fields with characteristic $2$ to general odd characteristics by applying the quadratic residue map $\eta$ and the theory of Kummer extensions of function fields. Since our construction is over elliptic function fields with odd characteristic, our constructions have flexible parameters. From the numerical experiments, it is noticed that the real correlation is much better than the theoretical upper bound given in Theorem III.6. In particular, for applications in cryptography, we also investigate the property of balance for this family of binary sequences. It turns out that the actual bound of $\Delta(\mathcal{S})$ is also much smaller than the theoretical bound. For the future work, it is interesting to 1) give better bounds for correlations and linear complexities of the sequences constructed in [9] and this paper by using other deep mathematical methods and 2) generalize the constructions of binary sequences by utilizing general algebraic function fields.

### v2

@@ I-A Our Main Results and Techniques
and correlation upper bounded by
$(2d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+2d$
where $\lfloor x\rfloor$ stands for the integer part of $x\in\mathbb{R}$. By taking explicit cyclic elliptic function fields, we extend the families of the binary sequences with new parameters, see Table I.
Now we overview the techniques which shall be used throughout this paper.
The automorphism group of an elliptic function field $E$ over $\mathbb{F}_{q}$ is in the form of the semi-product $T_{E}\rtimes\mathrm{Aut}(E,O)$, where $\mathrm{Aut}(E,O)$ consists of the $\mathbb{F}_{q}$-automorphisms of $E$ fixing the infinity place $O$, and $T_{E}$ corresponds to the translation group $E$. Let $Q$ be a place of $E$ with $\deg(Q)=d\geq 2$ such that $\gcd(d,q+1+t)=1$ and by the Riemann-Roch theorem, we have $\ell(Q)=\dim(\mathcal{L}(Q))=\deg(Q)+1-g(E)=\deg(Q)=d$ because of $\deg(Q)=d\geq 2g(E)-1=1$. Obviously, we have the decomposition $\mathcal{L}(Q)=\mathbb{F}_{q}\oplus V$. The vector space $V$ is called the complementary dual space of $\mathbb{F}_{q}$ in $\mathcal{L}(Q)$. The elements in $V\setminus\{0\}$ can be listed as $\{z_{1},z_{2},\cdots,z_{q^{d-1}-1}\}$.
@@
Then the family of binary sequences $\mathbf{s}_{i}$ has length $q+1+t$ and size $q^{d-1}-1$. The problems of computing the correlation are transformed into counting the number of rational places of certain Kummer extensions $E_{j,k}=E(y)$ where $y^{2}=z_{j}\cdot\tau^{-u}(z_{k})$ for $1\leq j,k\leq q^{d-1}-1,0\leq u\leq q+t$ by the cyclic structure of $T_{E}$.
@@
Although our bound applies to more general cases of the finite fields, both the zero places and pole places of $z_{j}\cdot\tau^{-u}(z_{k})$ are ramified in the Kummer extensions while only the pole places of $z_{j}+\tau^{-u}(z_{k})$ are ramified in the Artin-Schreier extensions for $1\leq j,k\leq q^{d-1}-1,0\leq u\leq q+t$, which induce the differences between the two bounds. Also by counting the rational places over the Kummer extensions, we also obtained the lower bound of the linear complexity.
On the other hand, we also explored the property of the balance. Also by transformation to the estimation of the rational places over the Kummer extensions, we have found that the theoretical value of the balance is upper bounded by $(d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+d.$ By utilizing the software SageMath [2], the actual difference or bound for correlation is much smaller than this theoretical bound as is shown in the tables in II-IV. As we have shown above, the main technique we utilize is the analysis for the behaviors of the rational places on the Kummer extensions over the cyclic elliptic function fields.
@@ I-B Organization of This Paper
The rest of this paper is organized as follows. In Section II, we provide preliminaries on binary sequences, extension theory of function fields, elliptic function fields and Kummer extensions of the function fields. In Section III, we present a theoretical construction of binary sequences with low correlation, favorable balance and large linear complexity via cyclic elliptic function fields over finite fields with odd characteristic by utilizing the theory of the Kummer extensions and we also present some examples to specify our construction. In Section IV, we give a conclusion and some possible future work on the constructions of binary sequences.
@@ III-B The Balance and Correlation
Now we define $\Delta(\mathcal{S})\triangleq\max\{|N^{\prime\prime}_{1}(\mathbf{s})-N^{\prime\prime}_{0}(\mathbf{s})|,\ \mathbf{s}\in\mathcal{S}\}$ and $\Delta(\mathcal{S})$ is called the balance for the family of sequences $\mathcal{S}$. In this subsection, we study the property of balance and it turns out that the problem is transformed into the analysis of the number of rational places for the Kummer extensions.
@@
$\Delta(\mathcal{S})\leq(d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+d.$ (2)
@@
On the other hand, we have $N(E^{\prime}_{i})=2N^{\prime\prime}_{0}(\mathbf{s}_{i})-|\mathcal{T}(z_{i})|$ for $1\leq i\leq q^{d-1}-1$ and $\mathcal{T}(z_{i})$ is denoted by the zeros that are rational of $z_{i}$. From the Serre bound, we have the estimation for the number of rational places over the curve $E^{\prime}_{i}$ and it is given by
@@
Let $n$ be an odd positive integer, $q=3^{n}$ and $t=\pm 3^{(n+1)/2}$. Then there exists a family of binary sequences with length $q+1+t$, size q1q-1 balance upper bounded by $\Delta(\mathcal{S})\leq 3\cdot\lfloor 2\cdot 3^{n/2}\rfloor+3^{(n+1)/2}+2$ and the correlation upper-bounded by $\mathrm{Cor}(\mathcal{S})\leq 5\cdot\lfloor 2\cdot 3^{n/2}\rfloor+3^{(n+1)/2}+4$.
@@
For $t=\pm 3^{(n+1)/2}$, there is a cyclic elliptic function field with $q+1+t$ rational places according to Lemma II.4. By Lemma II.5, there exists a place $Q$ in $E^{T_{E}}$ of degree $\deg(Q)=2$ that splits completely in $E$ as $\gcd(2,3^{n}+1+t)=1$. Thus, the corollary follows from Theorem III.6 and Theorem III.5. ∎
@@
Let $n$ be an odd positive integer, $q=3^{n}$ and t=1t=-1. Then there exists a family of binary sequences with length $q$, size q1q-1 balance upper bounded by $\Delta(\mathcal{S})\leq 3\cdot\lfloor 2\cdot 3^{n/2}\rfloor+2$ and the correlation upper-bounded by $\mathrm{Cor}(\mathcal{S})\leq 5\cdot\lfloor 2\cdot 3^{n/2}\rfloor+4$.
@@
For t=1t=-1, there is a cyclic elliptic function field with $q+1+t$ rational places according to Lemma II.4. By Lemma II.5, there exists a place $Q$ in $E^{T_{E}}$ of degree $\deg(Q)=2$ that splits completely in $E$ as $\gcd(2,3^{n})=1$. Thus, the corollary follows from Theorem III.6 and Theorem III.5. ∎
Corollary III.9.
Let $q=5^{n}$ for a positive integer $n$. Let $t=\pm 5^{n/2}$ and $n$ be even. Then there exists a family of binary sequences with length $q+1+t$, size q1q-1 balance upper bounded by $\Delta(\mathcal{S})\leq 3\cdot\lfloor 2\cdot 5^{n}\rfloor+|t|+2=7\cdot 5^{n/2}+2$ and correlation upper bounded by $\mathrm{Cor}(\mathcal{S})\leq 5\cdot\lfloor 2\cdot 5^{n}\rfloor+|t|+4=11\cdot 5^{n/2}+4$.
Proof.
For $t=\pm 5^{n/2}$, there is a cyclic elliptic function field with $q+1+t$ rational places according to Lemma II.4. By Lemma II.5, there exists a place $Q$ in $E^{T_{E}}$ of degree $\deg(Q)=2$ that splits completely in $E$ for the fact that $\gcd(2,5^{n}+1+t)=1$. Thus, the corollary follows from Theorem III.6 and Theorem III.5. ∎
@@ III-D Numerical Results
Field Size Length Family Size Bound Actual Balance
81 81 80 56 17
243 243 242 95 55
729 729 728 164 65
2187 2187 2186 281 97
Now we consider the actual correlation. Also let the field size be $q=3^{n}$. By the results in Theorem III.6, we can construct a family of binary sequences with size $3^{n}-1$, length $q$ and correlation upper bounded by $5\cdot\lfloor 2\sqrt{q}\rfloor+4$. We list some results calculated by SageMath [2] for this case in Table III.
@@
Field Size Length Family Size Bound Actual Correlation
@@
Field Size Sequence Length Family Size Bound Actual Correlation
@@ IV Conclusions and Future Work
In this paper, we extend the construction of binary sequences given by Jin et al. in [9] via the cyclic elliptic function fields with characteristic $2$ to general odd characteristics by applying the quadratic residue map $\eta$ and the theory of Kummer extensions of function fields. Since our construction is over elliptic function fields with odd characteristic, our constructions have flexible parameters. The balance between the numbers of 0s and $1$s, the correlation and the linear complexity of the binary sequences are investigated. Upper bounds for the balance and the correlation and a lower bound for the linear complexity are obtained. From the numerical experiments, it is noticed that the real value of balance or correlation is much better than the theoretical upper bound given in Theorem III.5 or Theorem III.6 respectively. For the future work, it is interesting to 1) give better bounds for correlations and linear complexities of the sequences constructed in [9] and this paper by using other deep mathematical methods and 2) generalize the constructions of binary sequences by utilizing general algebraic function fields.